OpenStudy (anonymous):

Which of the following options is an equivalent function to f(x) = 3(2)3x? f(x) = 3(8)x f(x) = 24x f(x) = 27(8)x f(x) = 3(8x) @nerlineg

OpenStudy (anonymous):

@nerlineg are you there?

OpenStudy (anonymous):

is the problem correct? (3)(2)(3x)?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

YES LOL IS IT RIGHT?

OpenStudy (astrophysics):

Are they all exponents?

OpenStudy (astrophysics):

You need to learn to put your answers right, if the function has exponents you put this little hat, 34x^2 like that, if it doesn't you put just x, we can't read minds.

OpenStudy (anonymous):

oh okay i know now for the next time

OpenStudy (astrophysics):

Ok so fix it now

OpenStudy (anonymous):

f(x) = 3(8)^x f(x) = 24^x f(x) = 27(8)^x f(x) = 3(8x)

OpenStudy (astrophysics):

And the original?

OpenStudy (anonymous):

yes

OpenStudy (astrophysics):

I mean, is what we're looking for related to f(x) = 3(2)3^x or is that just x..

OpenStudy (anonymous):

f(x) = 3(2)^3x

OpenStudy (astrophysics):

ok so its \[f(x) = 3(2)^{3x}\]

OpenStudy (anonymous):

ok so, thats the answer?

OpenStudy (astrophysics):

Well this isn't how it works, I will not give you the answer, because when it comes to your tests you will have no idea what you will be doing, like I said earlier you can graph these out yourself, but I know you won't do that. So I'll just leave you with this image and you can figure it out.  these are all the functions and this is the original, you can try figuring out what is going on here. 

OpenStudy (anonymous):

I DONT KNOW HOW?

OpenStudy (anonymous):

OH wait, the answer its f(x)= 3(8x)?

OpenStudy (astrophysics):

I think you should really review your exponent rules, because you don't even need graphs for this problem \[f(x) = 3(2)^{3x} \implies 3(8)^x\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

A?

OpenStudy (anonymous):

Identify the domain of the exponential function shown in the following graph:

OpenStudy (anonymous):

−5 ≤ x ≤ 5 0 ≤ x ≤ 5 All real numbers All positive numbers

OpenStudy (anonymous):

@nerlineg

OpenStudy (anonymous):

you know the answer?

OpenStudy (anonymous):

@ASAPT @chrisdbest @astrophysics

OpenStudy (anonymous):

@Vocaloid

OpenStudy (anonymous):

real numbers

OpenStudy (anonymous):

s not positive i see negatives

OpenStudy (anonymous):

thank you so much! so the answer its C?

OpenStudy (anonymous):

All real numbers?

OpenStudy (anonymous):

yes your right let me know if you got it right !

OpenStudy (anonymous):

yes it is thank you so much!! you're awesome

OpenStudy (anonymous):

Which inequality matches the graph?

OpenStudy (anonymous):

A)−2x + 3y > 7 B)2x − 3y < 7 C)−3x + 2y > 7 D)3x − 2y < 7

OpenStudy (anonymous):

@nerlineg

OpenStudy (anonymous):

if i was to guess ill pick B

OpenStudy (anonymous):

was that right?

OpenStudy (anonymous):

I THINK IS D

OpenStudy (anonymous):

I CHOOSE THAT

OpenStudy (anonymous):

did you get it right?